OpenStudy (anonymous):

The slope of the line below is -1. Write the point slope equation of the line using the coordinates of the labeled point (7,-7)

OpenStudy (igreen):

Point-Slope Form: \(\sf y - y_1 = m(x - x_1)\) Where \(\sf y_1\) is the y-value of the point, \(\sf x_1\) is the x-value of the point, and \(\sf m\) is the slope. Can you plug the numbers in?

OpenStudy (anonymous):

okay so i did y-y1=m(x-x1) y-(7)=(-1)(x-(-7) y-7=-1(x+7) y-7=-1x+(-7) y=-1x+-7-7 y=-1x-0 im not sure if i did this correctly

OpenStudy (anonymous):

is that right or no @iGreen

OpenStudy (igreen):

Not quite..you got 'x' and 'y' mixed up. (7, -7) x1 y1

OpenStudy (igreen):

Try again

OpenStudy (anonymous):

Okay so i redid it y-y1=m(x-x1) y-(-7)=(-1)(x-7) Y+7=-1(x-7) Y=7=-1x-(-7) y=-1x+7-7 Y=-1x+0 I got the same thing i think im doing something wrong @igreen

OpenStudy (igreen):

No..remember, they're only asking for the answer in point-slope form..there's no need to convert it to slope intercept form. y+7=-1(x-7) is correct

OpenStudy (anonymous):

ohhhh now i get it lol thank you im not the best at math lol

OpenStudy (anonymous):

can i ask you how to do one thing ?

OpenStudy (igreen):

Yes

OpenStudy (anonymous):

I have a question that asks how to find the point of a line described by an equation y-8=-3(x-6) how do i do that im not asking for the answer i wanna know how to do that

OpenStudy (igreen):

That equation is in point-slope form. And remember the definition of point-slope form. \(\sf y - y_1 = m(x - x_1)\) Where \(\sf y_1\) is the y-value of the point, \(\sf x_1\) is the x-value of the point, and \(\sf m\) is the slope. So what's the point in that equation?

OpenStudy (anonymous):

(6,8) ??

OpenStudy (anonymous):

and the slope is -3?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yep, you got it.